# CHIME-o-Grav: Wideband Timing of Four Millisecond Pulsars from the NANOGrav 15-yr dataset

### v2

@@ Abstract
Wideband timing of the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) datasets, where a single time-of-arrival (TOA) and a single dispersion measure (DM) are measured using the entire bandwidth of each observation, was first done for the 12.5-year dataset, and proved to be invaluable for characterizing the time-varying dispersion measure, reducing the data volume, and for improving the overall timing precision. The Canadian Hydrogen Intensity Mapping Experiment (CHIME) Telescope has been observing most NANOGrav millisecond pulsars (MSPs) at nearly daily cadence (compared to roughly monthly cadence for other NANOGrav observations) since 2019 with the objective of integration into future pulsar timing array (PTA) datasets. In this paper, we show the results of integration of high-cadence, low-observing-frequency CHIME data with data from the NANOGrav experiment for an isolated MSP PSR J0645++5158 and three binary MSPs PSR J1012++5307, PSR J2145-0750, and PSR J2302++4442. Using a wideband timing pipeline which we also describe, we present updated timing results for all four sources, including improvements in measurements of relativistic post-Keplerian parameters for the three binary pulsars in this analysis. For PSR J2302++4442, we report an updated strong detection of Shapiro delay from which we measured a companion mass of $0.35^{+0.05}_{-0.04}\ M_{\odot}$, a pulsar mass of $1.8^{+0.3}_{-0.3}\ M_{\odot}$, and an orbital inclination of ${80\degree}^{+1}_{-2}$. We also report updated constraints on the reflex motion for PSR J2145-0750 using a combination of Very Long Baseline Array astrometry and our updated measurement of the time derivative of the projected semi-major axis of the pulsar orbit as a prior.
[…]
1 Introduction
@@ 1.1 Wideband Timing
TOAs are usually calculated by cross-correlating a template of the integrated pulse profile to a noise-free template (Lorimer & Kramer, 2012). In traditional narrowband timing this template is frequency-averaged and does not account for evolution of the pulse profile across the observing bandwidth. When timing pulsars, we can expect some evolution of the pulse profile shape as a function of frequency caused by intrinsic changes in the pulsar magnetosphere, and extrinsic effects such as interstellar scattering and frequency-dependent dispersion measure (DM) (Hankins & Rickett, 1986; Alam et al., 2021a). Narrowband timing methods typically deal with profile evolution by calculating TOAs at several different frequencies in a single observation and the inclusion of arbitrary frequency dependent parameters in the timing model (Alam et al., 2021b).
@@
Wideband (WB) timing is a method of using a smooth frequency-dependent model, often called a template portrait, to simultaneously measure a TOA and a dispersion measure (DM) from each observation, without splitting the observations into multiple frequency sub-bands (Pennucci et al., 2014). The reference frequency for the TOA is selected such that there is zero covariance between the TOA phase offset and DM measurement. WB portraits are calculated from high SNR templates, referred to as data portraits, which are made by aligning and averaging many observations together (Pennucci, 2019).
@@
To construct the template portrait, principal component analysis (PCA) is first performed on the mean profile-subtracted data portrait. The resulting eigenvectors and the mean profile are then de-noised using a Stationary Wavelet Transform (Lee et al., 2019). The de-noised eigenvectors are then projected onto the mean profile to create coordinate curves representing the profile evolution which are approximated using spline interpolation. The coordinate splines and mean profile are used to reconstruct the 2-D template portrait which is then used for calculation WB TOAs.
@@
Calculating TOAs with this method eliminates some of the biases we see in narrowband timing and means that individual wideband TOAs have much greater precision than narrowband TOAs since only one TOA measurement need be made from a single observation. The use of wideband timing also significantly reduces the TOA volume in large PTA datasets, lowers the number of free parameters in timing models, and overall increases timing precision when compared to narrowband timing (Johnson et al., 2024; Agazie et al., 2023b; Alam et al., 2021a).
1.2 The Canadian Hydrogen Intensity Mapping Experiment
@@
The Canadian Hydrogen Intensity Mapping Experiment (CHIME) Telescope is a drift scan radio telescope that has been observing sources in the NANOGrav pulsar timing array with nearly daily cadence since 2019 (The CHIME Collaboration et al., 2022; Good, 2021). In this paper, we present results of wideband data combination of CHIME data with that from the NANOGrav 15 year dataset (NG15) for four pulsars: PSR J0645++5158, PSR J1012++5307, PSR J2145-0750, and PSR J2302++4442. In Section 2 we describe the instruments used for observing and the data reduction techniques used to calculate the CHIME WB TOAs. Section 3 describes the binary models, DM modeling, and noise modeling used for the timing analysis of the combined CHIME-NG15 dataset. Section 4 reports timing analysis results for each source, and Section 5 discusses the impact of the inclusion of CHIME data of our timing results. Finally, Section 6 summarizes our major conclusions and areas for future work.
2 Observations and Data Reduction
@@ 2.3 Data Reduction
For CHIME data, we developed a data processing pipeline chimera to perform equivalent processing steps to those in nanopipe specifically for generating WB TOAs. This pipeline is described in Appendix A. The exception being that we did not do polarization calibration of the CHIME data, as calibration methods for CHIME are still under development (CHIME/Pulsar Collaboration et al., 2021).
2.4 WB Portraits
Generation of WB templates and TOA calculation was done using the PulsePoratraiture [^3] https://github.com/pennucci/PulsePortraiture software suite (Pennucci, 2019). WB templates are made by first constructing a high S/N data portrait by adding and aligning all observations from a given receiver together using ppalign.py. From these data portraits, the model portraits are calculated using a combination of principal component analysis and spline interpolation using ppspline.py. The choice of the number of significant eigenvectors was done using an S/N threshold of 150, which was calculated using methods from Pennucci (2019) and NANOGrav Collaboration et al. (2015).
WB model portraits do not inherently account for instrument offsets and are not accurate beyond the frequency range of the data portrait template (Pennucci, 2019). The WB templates used for NG15 data were the same as those used in Alam et al. (2021a). We required the calculation of new CHIME-specific portraits for each source. To avoid the inclusion of noisy or marginal detections, only observations with an S/N exceeding an empirically determined threshold, we chose the 50th percentile of S/N per pulsar, were included in the data portrait. This threshold was 60 for PSR J1012++5307 and 20 for PSR J0645++5158, J2145-0750, and J2302++4442.
@@ 3.3 Noise Modeling
Where $A_{\rm red}$ is the power spectrum amplitude at a frequency of 1 yr-1 and $\gamma_{\rm red}$ is the spectral index. Similarly to Alam et al. (2021a) we considered detection of red noise to be significant if the Savage-Dickey density ratio of the posterior distribution of $\log_{10}(A_{\rm red})$ exceeded 100 (Dickey, 1971). We detected significant red noise in PSR J1012++5307, which was also detected in NG15. We get somewhat different values for $\gamma_{\rm red}$, but the NG15 value of 0.8(3)-0.8(3) is consistent within the 2-$\sigma$ confidence interval of the CHIME-NG15 result of 1.3(4)-1.3(4). For PSR J2145-0750, there was significant red noise in the NG15 data, but with the combined CHIME-NG15 dataset our attempts to model red noise would introduce large, non-physical sinusoids in our timing residuals so we excluded it from the noise model and only used white noise parameters. We believe this difficultly may be due to the strong solar wind variability we see the DM times series (Section 3.2).
4 Results
4.1 PSR J0645++5158
@@ 4.2 PSR J1012++5307
The uncertainty of $\dot{P}_{B}^{\rm Shk}$ is dominated by the timing parallax measurement which, for PSR J1012++5307, was consistent within 1.4-$\sigma$ to Very Long Baseline Interferometry (VLBI) determined parallax measurements from Ding et al. (2023) (Table 4). The larger uncertainty in the CHIME-NG15 timing parallax vs VLBI parallax explains the larger uncertainty of our estimate of $\dot{P}_{B}^{\rm Shk}$, though still consistent within 1.1-$\sigma$ of the $\dot{P}_{B}^{\rm Shk}$ estimate from VLBI astrometry (Table 5).
@@
The measurement of rate of change of the projected semi-major axis ($\dot{x}$) for PSR J1012++5307 improved in precision by a factor of three from the NG15 value. Using our own measurement uncertainty, the NG15 value is 5 standard deviations from our measurement, but when using the larger NG15 measurement uncertainty, our measurement was within two standard deviations. Contributors to the measured value of $\dot{x}$ are summarized below:
@@ 4.4 PSR J2302++4442
Comparing this result to the NG15 Shapiro delay results, we switched the binary model from DD to ELL1H due to the nearly circular eccentricity, and thus used a different parameterization. In the DD binary model, Shapiro delay is parameterized in the timing model by $m_{\rm c}$ and $\sin i$ (Damour & Deruelle, 1985, 1986). The combined CHIME-NG15 results show significant improvement from the Shapiro-delay estimated mass measurements from Agazie et al. (2023b), which reported an $m_{\rm c}$ of $0.3\pm 0.1\,M_{\odot}$, $m_{\rm p}$ of $5.2^{+3.0}_{-3.2}\,M_{\odot}$, and system inclination of $78\degree\pm 3.3\degree$. Our results for $m_{\rm c}$ and $i$ have improved precision by a factor of 2 and are consistent within the 1-$\sigma$ error region. Our $m_{\rm p}$ measurement uncertainty is dramatically lower by two orders of magnitude and is well-determined in a more physically motivated range of pulsar masses than the analysis in Agazie et al. (2023b). This improvement is possibly due to a combination of the longer timing baseline having more full orbits, this pulsar has a long binary period of 125 days, and the change to the ELL1H binary model to account for the nearly circular orbit and high system inclination.
Figure 6: PSR J2302++4442 measurement of companion mass and cosine of inclination ($\cos i$ from Shapiro delay. Center plot is the 2-D posterior pdf in purple for $m_{\rm c}$ and $\cos i$ with 1,2,3-$\sigma$ contours overlaid. The grey cross in the center is the analytically determined 68% confidence intervals for $m_{\rm c}$ and $\cos i$. Top plot is the marginalized 1-D PDF for $m_{c}$, and right side plot is the marginalized 1-D PDF for $\cos i$. Solid red lines indicate the measurement value, and dotted red lines indicate bounds of the 68% confidence intervals.
@@
4.5 Solar wind modeling
@@
5 Discussion
The combination of four years of CHIME data with NG15 data have already massively increased the TOA volume per source. Even for sources like PSR J1012++5307 and PSR J2145-0750 that have the longest NG15 timing baseline, there are already enough CHIME TOAs to nearly triple the total TOA volume (Table 1). This increase will get compounded in narrowband datasets made with CHIME data as there could be anywhere from several to dozens of TOAs per observations to allow proper frequency coverage. As CHIME data is currently being incorporated into multiple upcoming PTA datasets, such as NANOGrav’s next dataset and IPTA’s upcoming third data release, this sudden explosion of TOAs is likely to create a data volume issue for both single source analyses and PTA analyses.
5.1 PK parameter improvements
The fractional uncertainty of PK parameters is expected to decrease with increasing observing baseline. The $\dot{P}_{B}$ measurement for PSR J1012++5307 decreased in fractional uncertainty between the 12.5 yr dataset and NG15 by the theoretical expected improvement of $T^{-5/2}$ (Damour & Taylor, 1992; Alam et al., 2021a; Agazie et al., 2023b), where $T$ represents the observing time in years. Similarly the improvement in fractional uncertainty with the inclusion of four years of CHIME data to NG15 was also $T^{-5/2}$. We also calculated that the fractional uncertainty of $\dot{x}$ for PSR J1012++5307 improved by roughly $T^{-2.1}$ between NG15 and our CHIME-NG15 analysis. This parameter was newly detected for this source NG15 so we could not compare to the 12.5 yr dataset. We were able to do a comparison for the $\dot{x}$ measurement for PSR J2145-0750, where the 12.5-yr to NG15 improvement was $T^{-2.3}$ and the NG15 to CHIME-NG15 improvement was $T^{-2.2}$. These very similar improvements in fractional uncertainty likely indicate that the higher cadence of the CHIME data is less important for the improvement of PK parameter measurements. The inclusion of CHIME data mainly benefits these PK measurements by significantly increasing the timing baseline.
5.2 CHIME Frequency coverage and Wideband TOA precision

### v3

@@ Abstract
Wideband timing of the the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) datasets was first done for the 12.5 yr dataset. This method, where a single time-of-arrival (TOA) and a single dispersion measure (DM) are measured using the entire bandwidth of each observation, proved to be invaluable for characterizing the time-varying dispersion measure, improving handling of frequency dependent profile variability, and data volume reduction. The Canadian Hydrogen Intensity Mapping Experiment (CHIME) Telescope has been observing most NANOGrav millisecond pulsars (MSPs) at nearly daily cadence (compared to roughly monthly cadence for other NANOGrav observations) since 2019 with the objective of integration into future pulsar timing array (PTA) datasets. In this paper, we show the results of integration of high-cadence, low-observing-frequency CHIME data with data from the NANOGrav experiment for an isolated MSP PSR J0645++5158 and three binary MSPs PSR J1012++5307, PSR J2145-0750, and PSR J2302++4442. Using a wideband timing pipeline which we also describe, we present updated timing results for all four sources, including improved relativistic post-Keplerian measurements for the three binary pulsars in this analysis. For PSR J2302++4442, we report an updated strong detection of Shapiro delay from which we measured a companion mass of $0.35^{+0.05}_{-0.04}\ M_{\odot}$, a pulsar mass of $1.8^{+0.3}_{-0.3}\ M_{\odot}$, and an orbital inclination of ${80\degree}^{+1}_{-2}$. We also report updated constraints on the reflex motion for PSR J2145-0750 using a combination of Very Long Baseline Array astrometry and our updated measurement of the time derivative of the projected semi-major axis of the pulsar orbit as a prior.
I Introduction
@@ I.1 Wideband Timing
TOAs are usually calculated by cross-correlating observed pulse profiles with a noise-free template of the integrated pulse profile (Lorimer & Kramer, 2012). In traditional narrowband timing this template is frequency-averaged and does not account for evolution of the pulse profile across the observing bandwidth. When timing pulsars, we can expect some evolution of the pulse profile shape as a function of frequency caused by intrinsic changes in the pulsar magnetosphere, and extrinsic effects such as interstellar scattering and frequency-dependent dispersion measure (DM) (Hankins & Rickett, 1986; Alam et al., 2021a). Narrowband timing methods typically deal with profile evolution by calculating TOAs at several different frequencies in a single observation and the inclusion of arbitrary frequency dependent parameters in the timing model (Alam et al., 2021b).
@@
Wideband (WB) timing is a method of using a smooth 2-D frequency-dependent model of the pulse profile, often called a template portrait, to simultaneously measure a TOA and a dispersion measure (DM) from each observation, without splitting the observations into multiple frequency sub-bands (Pennucci et al., 2014). The reference frequency for the TOA is selected such that there is zero covariance between the TOA phase offset and DM measurement. WB template portraits are calculated from high signal to noise (S/N) data portraits, which are constructed by aligning and averaging many observations with strong detections of the pulsar together (Pennucci, 2019).
@@
To construct the template portrait, the mean (frequency-averaged) profile-subtracted data portrait is decomposed into an orthonormal set of eigenprofiles using principal component analysis (PCA). A small number of the eigenprofiles are kept as the basis for constructing the profile evolution model. The eigenprofiles and the mean profile are then de-noised using a Stationary Wavelet Transform (Lee et al., 2019). Coordinate functions of the profile evolution are constructed by projecting the de-noised eigenprofiles onto the de-noised mean profile. These are then approximated by fitting a spline to each coordinate function. The coordinate function splines and de-noised mean profile are used to construct the 2-D template portrait which is then used for calculation WB TOAs.
@@
Calculating TOAs with this method means that individual wideband TOAs often have much greater precision than individual narrowband TOAs derived from the same data. Only one wideband TOA measurement need be made from a single observation to cover the total observing bandwidth, whereas multiple narrowband TOAs are needed to cover the total bandwidth and then are fit simultaneously, with a final precision comparable to that of the wideband data. However, using multiple narrowband TOAs may require arbitrary frequency dependent parameters in the timing model because profile evolution is handled at the timing stage rather than the profile template creation stage. The use of wideband timing significantly reduces the TOA volume in large PTA datasets, lowers the number of free parameters in the timing model, and more smoothly handles frequency dependent profile variability when compared to narrowband timing (Johnson et al., 2024; Agazie et al., 2023b; Alam et al., 2021a).
I.2 The Canadian Hydrogen Intensity Mapping Experiment
@@
The Canadian Hydrogen Intensity Mapping Experiment (CHIME) Telescope is a drift scan radio telescope that has been observing sources in the NANOGrav pulsar timing array with nearly daily cadence since 2019 (The CHIME Collaboration et al., 2022; Good, 2021). In this paper, we present results of wideband data combination of CHIME data with that from the NANOGrav 15 year dataset (NG15) for four pulsars: PSR J0645++5158, PSR J1012++5307, PSR J2145-0750, and PSR J2302++4442. In Section II we describe the instruments used for observing and the data reduction techniques used to calculate the CHIME WB TOAs. Section III describes the binary models, DM modeling, and noise modeling used for the timing analysis of the combined CHIME/NG15 dataset. Section IV reports timing analysis results for each source, and Section V discusses the impact of the inclusion of CHIME data of our timing results. Finally, Section VI summarizes our major conclusions and areas for future work.
II Observations and Data Reduction
@@ II.3 Data Reduction
For CHIME data, we developed a data processing pipeline chimera (Susobhanan, 2026) to perform equivalent processing steps to those in nanopipe specifically for generating WB TOAs. This pipeline is described in Appendix A. The exception being that we did not do polarization calibration of the CHIME data, as calibration methods for CHIME are still under development (CHIME/Pulsar Collaboration et al., 2021). Polarization calibration would account for instrumental polarization distortions which could introduce systemic errors into TOA calculations if not corrected for (van Straten, 2026). These errors can be particularly pronounced for sources with a high polarization fraction as the pulsar signal passing through the interstellar medium can also undergo significant Faraday rotation. Of the sources in this analysis, PSR J1012++5307 and PSR J2145-0750 have been previously observed to have significant polarization fractions of 0.66 and 0.21 respectively (Wahl et al., 2022). Introducing polarization calibration in future analyses could account for instrumental errors, reduce TOA uncertainty, and improve timing rms residuals.
II.4 WB Template Portraits
Generation of WB template portraits and TOA calculation was done using the PulsePoratraiture [^3] https://github.com/pennucci/PulsePortraiture software suite (Pennucci, 2019). WB template portraits are made by first constructing a high S/N data portrait by adding and aligning all observations from a given receiver together using ppalign.py. From these data portraits, the template portraits are calculated using a combination of principal component analysis and spline interpolation using ppspline.py. The choice of the number of significant eigenprofiles was done using an S/N threshold of 150, which was calculated using methods from Pennucci (2019) and NANOGrav Collaboration et al. (2015).
WB template portraits do not inherently account for instrument offsets and are not accurate beyond the observing frequency range of the data portrait (Pennucci, 2019). The WB template portraits used for NG15 data were the same as those used in Alam et al. (2021a). We required the calculation of new CHIME-specific template portraits for each source. To avoid the inclusion of noisy or marginal detections, only observations with an S/N exceeding an empirically determined threshold, we chose the 50th percentile of S/N per pulsar, were included in the construction of the data portraits. This threshold was 60 for PSR J1012++5307 and 20 for PSR J0645++5158, J2145-0750, and J2302++4442.
@@ III.3 Noise Modeling
Where $A_{\rm red}$ is the power spectrum amplitude at a frequency of 1 yr-1, f is the Fourier frequency, and $\gamma_{\rm red}$ is the spectral index. Similarly to Alam et al. (2021a) we considered detection of red noise to be significant if the Savage-Dickey density ratio of the posterior distribution of $\log_{10}(A_{\rm red})$ exceeded 100 (Dickey, 1971). We detected significant red noise in PSR J1012++5307, which was also detected in NG15. We get somewhat different values for $\gamma_{\rm red}$, but the NG15 value of 0.8(3)-0.8(3) is consistent within the 2-$\sigma$ confidence interval of the CHIME/NG15 result of 1.3(4)-1.3(4). The Fourier frequencies used to fit for $A_{\rm red}$ and $\gamma_{\rm red}$ are dependent on the total timing baseline, which is longer by nearly three years for CHIME/NG15, thus we fitting these parameters using somewhat different frequencies. Additionally, the added CHIME data has both a higher observing cadence and increased TOA uncertainties, which implicitly changes the weighting of the underlying basis functions used for fitting. Both PSR J0645++5158 and PSR J2302++4442 did not have significant detections of red noise in both this analysis and NG15. For PSR J2145-0750, there was significant red noise in the NG15 data, but with the combined CHIME/NG15 dataset our attempts to model red noise would introduce large, non-physical sinusoids in our timing residuals so we excluded it from the noise model and only used white noise parameters. We believe this difficultly may be due to the strong solar wind variability we see the DM times series (Sections III.2 & IV.5). If we had been able to model red noise we would expect to see some increases in timing parameter uncertainty and possibly some changes in post-keplerian parameter measurements as unmodeled red noise processes would vary on similar timescales to relativistic post-keplerian processes. For PSR J1012++5307 we did observe some changes to post-keplerian parameters orbital decay and rate of change of the projected semi-major axis prior to and after introducing a red noise model that were within 2-$\sigma$ measurement uncertainties.
IV Results
IV.1 PSR J0645++5158
@@ IV.2 PSR J1012++5307
The uncertainty of $\dot{P}_{B}^{\rm Shk}$ is dominated by the timing parallax measurement which, for PSR J1012++5307, was consistent within 1.4-$\sigma$ to Very Long Baseline Interferometry (VLBI) determined parallax measurements from Ding et al. (2023) (Table 4). The larger uncertainty in the CHIME/NG15 timing parallax vs VLBI parallax explains the larger uncertainty of our estimate of $\dot{P}_{B}^{\rm Shk}$, though still consistent within 1.1-$\sigma$ of the $\dot{P}_{B}^{\rm Shk}$ estimate from VLBI astrometry (Table 5).
@@
The measurement of rate of change of the projected semi-major axis ($\dot{x}$) for PSR J1012++5307 improved in precision by a factor of three from the NG15 value. Using our own measurement uncertainty, the NG15 value is 5 standard deviations from our measurement, suggesting the values may be significantly different, but when using the larger NG15 measurement uncertainty, our measurement was within two standard deviations. Thus we can treat these measurements as broadly consistent, with the CHIME/NG15 measurement having improved precision due to the longer timing baseline. Contributors to the measured value of $\dot{x}$ are summarized below:
@@ IV.4 PSR J2302++4442
Comparing this result to the NG15 Shapiro delay results, we switched the binary model from DD to ELL1H due to the nearly circular eccentricity, and thus used a different parameterization. In the DD binary model, Shapiro delay is parameterized in the timing model by $m_{\rm c}$ and $\sin i$ (Damour & Deruelle, 1985, 1986). The combined CHIME/NG15 results show significant improvement from the Shapiro-delay estimated mass measurements from Agazie et al. (2023b), which reported an $m_{\rm c}$ of $0.3\pm 0.1\,M_{\odot}$, $m_{\rm p}$ of $5.2^{+3.0}_{-3.2}\,M_{\odot}$, and system inclination of $78\degree\pm 3.3\degree$. Our results for $m_{\rm c}$ and $i$ have improved precision by a factor of 2 and are consistent within the 1-$\sigma$ error region. Our $m_{\rm p}$ measurement uncertainty is dramatically lower by two orders of magnitude and is well-determined in a more physically motivated range of pulsar masses than the analysis in Agazie et al. (2023b). This improvement is possibly due to a combination of the longer timing baseline having more full orbits, this pulsar has a long binary period of 125 days, and the change to the ELL1H binary model to account for the nearly circular orbit and high system inclination.
Figure 6: PSR J2302++4442 measurement of companion mass and cosine of inclination ($\cos i$) from Shapiro delay. Center plot is the 2-D posterior pdf in purple for $m_{\rm c}$ and $\cos i$ with 1,2,3-$\sigma$ contours overlaid. The grey cross in the center is the analytically determined 68% confidence intervals for $m_{\rm c}$ and $\cos i$. Top plot is the marginalized 1-D PDF for $m_{c}$, and right side plot is the marginalized 1-D PDF for $\cos i$. Solid red lines indicate the measurement value, and dotted red lines indicate bounds of the 68% confidence intervals.
@@
IV.5 Solar wind modeling
@@
V Discussion
The combination of four years of CHIME data with NG15 data have already massively increased the TOA volume per source. Even for sources like PSR J1012++5307 and PSR J2145-0750 that have the longest NG15 timing baseline, there are already enough CHIME TOAs to nearly triple the total TOA volume (Table 1). This increase will get compounded in narrowband datasets made with CHIME data as there could be anywhere from several to dozens of TOAs per observations to allow proper frequency coverage. For comparison, in NG15 the narrowband datasets for the sources in this paper had at least 41 times the number of TOAs as the corresponding NG15 WB datasets. As CHIME data is currently being incorporated into multiple upcoming PTA datasets, such as NANOGrav’s upcoming 20-year dataset and IPTA’s third data release, this sudden explosion of TOAs is likely to create a data volume issue for both single source analyses and PTA analyses.
V.1 PK parameter improvements
The fractional uncertainty of PK parameters is expected to decrease with increasing observing baseline. The $\dot{P}_{B}$ measurement for PSR J1012++5307 decreased in fractional uncertainty between the 12.5 yr dataset and NG15 by the theoretical expected improvement of $T^{-5/2}$ (Damour & Taylor, 1992; Alam et al., 2021a; Agazie et al., 2023b), where $T$ represents the observing time in years. Similarly the improvement in fractional uncertainty with the inclusion of four years of CHIME data to NG15 was also $T^{-5/2}$. We also calculated that the fractional uncertainty of $\dot{x}$ for PSR J1012++5307 improved by roughly $T^{-2.1}$ between NG15 and our CHIME/NG15 analysis. This parameter was newly detected for this source NG15 so we could not compare to the 12.5 yr dataset. We were able to do a comparison for the $\dot{x}$ measurement for PSR J2145-0750, where the 12.5-yr to NG15 improvement was $T^{-2.3}$ and the NG15 to CHIME/NG15 improvement was $T^{-2.2}$. These very similar improvements in fractional uncertainty likely indicate that the higher cadence of the CHIME data is less important for the improvement of PK parameter measurements. The inclusion of CHIME data mainly benefits these PK measurements by significantly increasing the timing baseline.
V.2 CHIME Frequency coverage and WB TOA precision